Tech Staff Meeting Summary 2/8/02

CVTN is running. Need one more visit to get comm going.

CBHT comm. Switch combined with dramatic k2ew config changes appeared to work. Steve got said a new firmware upgrade may also solve the problem.

HDBT. Need a fence and gravel for ground cover.

NAIT. NWS does not know when the VPN will be installed.

CUET. Still waiting to hear back from Pat Bohannan at TEMA regarding firewall holes. Comm should be ready to come up as soon as Nashville allows external access.

SFTN. K2 and MSS100 are fine. No comm out the ethernet. Needs debugging.

TUMT. Occasionally loses comm. Likely config issue with k2ew.

DBB. Steve suggested a terminal server/watchdog to do dynamic resets when the hang. He will do a mockup to test the idea. Mitch will look into the ISTI worm module to monitor outages. Jim will check into getting the dish hung at NMAD.

Jim got some additional quotes for the NMAD fence. Need to make requisition.

GPS. Still no detail in GPS status. Can't fix what we don't know is broke.

Still working on the box for the sensors in the yard. Steve got the 4 port terminal server. Mitch will configure. Need the rest of the guts.

ET. John has been spending alot of time on PII boxes, destroyed and moved stations, etc. Need to discuss when John gets here. Need to closely look at what is really needed in NC. Mitch is rerouting the internet communications to try to reduce the load. Need to make a trip over there to help get things running. Need a couple aironets at ATHN to get it online.

NM. There are 5 New Madrid stations to fix yet, and a bunch of little things.

Chapman and Munsey are working on a site for Oakridge to put a CMG3 or 40T.

FDL. Russ has the two broken K2's. Will use aftershock params from Paul as "standard" config to say they are ready to go. Any customization will need to be done just prior to deployment. Also need to do "end to end" tests.

PII. Most of PI has been cleaned out. Still need to go through PII but it is in relatively good shape. Need to get going on tuning and calibrating.

Monroe. The current site is under water. Don Wheeler is consulting with Greg to find a new site.

RPC. Mitch has acquired funds for the Regional Processing Center hardware (RPC). Greg is going to order the racks. Have two 48VDC supplies. Still need the rest of the power system, but there is enough for Mitch and Steve to plug them in and being developing configuration.

FY02 ANSS. Will be bringing techs from USC, SLU, VPI, and Parker to Memphis. Don't know what we're installing so we must assume it won't be too much different from last year.

Greg is working on version 95g of the I-40 installation.

SFTN needs solar backup power.

Need to do site visits of Lamont stations sometime this year.